A 1000kg car moving at 97.2 kph is stopped in 2 seconds, how much force was required to this?


1
Expert's answer
2022-01-12T08:28:23-0500

The initial momentum of the car is:


"p_i = 1000kg\cdot 97.2kph = 1000kg\cdot 27m/s = 27000kg\cdot m/s"

The final momentum of the car is "p_f = 0". Thus, the force reqired to stop can be found from the second Newton's law:


"F = \dfrac{p_i-p_f}{\Delta t}"

where "\Delta t = 2s". Thus, obtain:


"F = \dfrac{27000kg\cdot m/s-0}{2s} = 13500N"

Answer. 13500 N.